### v2

# Handling Domain Shifts for Anomalous Sound Detection: A Review of DCASE-Related Work

Abstract
When detecting anomalous sounds in complex environments, one of the main difficulties is that trained models must be sensitive to subtle differences in monitored target signals, while many practical applications also require them to be insensitive to changes in acoustic domains. Examples of such domain shifts include changing the type of microphone or the location of acoustic sensors, which can have a much stronger impact on the acoustic signal than subtle anomalies themselves. Moreover, users typically aim to train a model only on source domain data, which they may have a relatively large collection of, and they hope that such a trained model will be able to generalize well to an unseen target domain by providing only a minimal number of samples to characterize the acoustic signals in that domain. In this work, we review and discuss recent publications focusing on this domain generalization problem for anomalous sound detection in the context of the DCASE challenges on acoustic machine condition monitoring.
Index Terms:
anomalous sound detection, domain shift, domain adaptation, domain generalization, machine condition monitoring
1 Introduction
[…]
asd has a wide range of applications, including acoustic monitoring of machines [1, 2, 3, 4, 5], health [6, 7], roads [8], smart home environments [9], or public places [10]. The goal in all these applications is to distinguish between normal and anomalous audio recordings. Usually, anomalous sound detection (ASD) systems are trained exclusively with normal data, as anomalous data is often difficult and costly to obtain.
One of the major difficulties that ASD systems need to overcome is how to robustly handle so-called domain shifts, changes in the recorded audio signal that are caused by changes in acoustic environments, sensors, or properties of monitored sound sources themselves. Inherently, domain shifts have a strong impact on the audio signal and thus also affect the outcome of ASD systems if no precautions are taken. However, in most applications, this effect is not desirable and should be suppressed. Ideally, trained ASD systems should be completely insensitive to domain shifts, while being very sensitive to modifications of the monitored target signals that indicate the occurrence of application-dependent anomalies. As these differences may be very subtle compared to changes caused by domain shifts, especially in noisy environments with many sound sources, developing ASD systems that perform well in domain-shifted conditions is very challenging.
This paper, which extends a non-peer reviewed manuscript [11], reviews the latest works on how to handle domain shifts for ASD in the context of the annual DCASE challenge [12]. We first describe the challenges posed by domain shifts and the resulting domain mismatch, illustrated in Fig. 1. We then describe the two main types of approaches to handling them, domain adaptation (DA) and domain generalization (DG), defining the terms, listing publicly available datasets, and reviewing recently published works related to these topics. Last but not least, the limitations of different approaches are discussed and future directions of research are identified.
Figure 1: Illustration of the domain mismatch between the anomaly scores of a source domain and a target domain. Normal and anomalous samples are usually less well separated in the target domain than in the source domain, which decreases domain-independent performance over the performance obtained for the source domain alone. Furthermore, the optimal decision thresholds for separating the scores belonging to the normal and anomalous data of different data domains differ substantially, which significantly decreases performance when using only a single threshold for both domains.
Table 1: DCASE ASD datasets in domain-shifted conditions.
# recordings (per section) # machine types # sections (per machine type) source domain target domain Name Purpose total dev. set eval. set total dev. set eval. set split normal anomalous normal anomalous supplemental DCASE2021 [1] DA $7$ 7 7 6 3 3 train $1000$ 00 $3$ 00 00 DCASE2021 [1] DA $7$ 7 7 6 3 3 test $100$ $100$ $100$ $100$ 00 DCASE2022 [2] DG $7$ 7 7 6 3 3 train $990$ 00 $10$ 00 00 DCASE2022 [2] DG $7$ 7 7 6 3 3 test $50$ $50$ $50$ $50$ 00 DCASE2023 [3] DG $14$ 7 7 1 1 1 train $990$ 00 $10$ 00 00 DCASE2023 [3] DG $14$ 7 7 1 1 1 test $50$ $50$ $50$ $50$ 00 DCASE2024 [4] DG $16$ 7 9 1 1 1 train $990$ 00 $10$ 00 00 DCASE2024 [4] DG $16$ 7 9 1 1 1 test $50$ $50$ $50$ $50$ 00 DCASE2025 [5] DG $14$ 7 7 1 1 1 train $990$ 00 $10$ 00 $100$ DCASE2025 [5] DG $14$ 7 7 1 1 1 test $50$ $50$ $50$ $50$ 00
2 Domain Shifts
The goal of ASD is to determine, based on a sound recording of a phenomenon, whether that phenomenon is normal or anomalous. Its normal/anomalous characteristic is an intrinsic property of that phenomenon: as such, the determination by the ASD system should ideally be independent of the conditions in which the sound was recorded, as long as the change in conditions does not alter the observability and nature of the phenomenon’s normal/anomalous characteristic. Such changes include the use of microphones of different types or in different locations, the presence or absence of other sound sources, or modifications in certain properties of the monitored sound sources themselves, e.g., the use of different settings of the monitored machines.
ASD systems are however often placed in a conundrum: their only view of the phenomenon’s normal/anomalous characteristic is via a training set of normal data recorded under a certain set of conditions, referred to as a source domain, and they are expected to make a determination on the normal/anomalous characteristic on a new sample recorded under a potentially different set of conditions, referred to as a target domain, for which they only have a few training samples. Changes in the recording conditions occurring between the source domain and the target domain constitute a so-called domain shift. ASD systems thus need to implicitly disentangle the effects on the sound signal of the normal/anomalous characteristic and the recording conditions, in order to be robust to changes in these conditions.
An additional difficulty lies in the large imbalance often observed in practice between the number of available training samples in the source and target domains. Many training samples are typically available for the source domain, while only a few training samples may be available for the target domain. Recording a large amount of new training samples after a domain shift occurred may be impractical, as it requires much effort, and thus is very costly; domain shifts may also frequently or even continuously occur; and operators may not even be aware of such domain shifts.
A consequence of domain shifts related to differences in signal space is that anomaly scores are usually also distributed very differently across both domains. This results in a so-called domain mismatch, illustrated in Fig. 1. Since the embedding models were trained with no or few samples belonging to the target domain, the anomaly score distributions for the normal and anomalous samples are not well-separated in the target domain, which inherently leads to a worse ASD performance than in the source domain. Furthermore, the optimal decision thresholds differ substantially, further degrading the performance if the same decision threshold is used for both domains.
3 Domain Adaptation
One way to handle domain shifts is to adapt an existing ASD system that was trained with sufficient data from a source domain to a specific target domain. Here, the main challenge is that the training set for the target domain consists of only very few samples and thus knowledge from the source domain, which may differ substantially from the target domain, needs to be transferred somehow to obtain a well-performing system. Note that once a system is adapted to a target domain, it may no longer perform well in the source domain it was trained on.
3.1 Datasets
[…]
3.2 Approaches
[…]
Topic Strategy Reference Main idea DCASE datasets used DA - Chen et al [15] fine-tune model 2021 DA - Kuroyanagi et al [16] fine-tune models and domain-specific backends 2021 DA - Yamaguchi et al [17] fine-tune batch normalization layers other DA - Lopez et al [18] domain-specific normalization with AutoDIAL 2021 DA - Chen et al [19] gradient-based meta learning 2021 DA - Wilkinghoff [20] domain-specific backends 2021 DG domain specialization Kuroyanagi et al [21] domain-specific models and domain classifier 2022 DG domain specialization Kuroyanagi et al [21] balance mini-batches with mixup 2022 DG domain specialization Guan et al [22] balance domains with SMOTE 2022 DG domain specialization Junjie et al [23] balance domains with SMOTE 2023 DG domain specialization Chen et al [24] balance domains with SMOTE and a mixup variant 2022,2023 DG domain-invariant representations Deng et al [25] normalize samples independently for each batch 2022 DG domain-invariant representations Nejjar et al [26] VICreg extension with mixup 2022 DG domain-invariant representations Yan et al [27] minimize covariance differences between domains 2020, 2022 DG feature disentanglement Venkatesh et al [28] discriminative tasks for sections and attributes 2022 DG feature disentanglement Dohi et al [29] disentangle attributes with normalizing flow other DG feature disentanglement Lan et al [30] Mahalanobis distances for hierarchical metadata 2022 DG feature disentanglement Guan et al [31] gradient reversal with hierarchical metadata 2022 DG anomaly score calculation Harada et al [32] domain-specific Mahalanobis distance 2022 DG anomaly score calculation Wilkinghoff [33] nearest neighbor based anomaly scores 2022 DG anomaly score calculation Wilkinghoff et al [34] local density based score normalization 2020, 2023, 2024 DG anomaly score calculation Saengthong et al [35] domain-wise standardization of scores 2020, 2023
Approaches for DA mostly focus on first training an ASD system with data from the source domain and then adapting this trained system to a specific target domain. One possibility to do this is to fine-tune an entire model that was trained in the source domain with data belonging to the target domain [16, 15]. This changes the problem of balancing the very differently sized training datasets of both domains to a problem of preventing the model from overfitting to the few target domain samples available for training. In [17], only the parameters of batch normalization layers [36] are fine-tuned to the target domain to minimize the computational costs needed for the adaptation while also reducing overfitting effects. A similar idea based on domain-specific normalization layers is realized in [18] by using AutoDIAL [37]. The authors of [19] propose to use gradient-based meta learning [38] and a prototypical loss [39] in order to more effectively adapt to target domains with only a very small amount of training samples. Another DA approach is to simply train a joint embedding model for both domains but estimate the distributions of each domain individually [20, 16]. Table 2 briefly summarizes all mentioned approaches.
4 Domain Generalization
Adapting models for each unknown domain with the possible need of re-training models, fine-tuning hyperparameters, or even replacing system components is very impractical, as it is computationally costly and may require expert knowledge. It is thus much more desirable to strive for domain generalization (DG) [40, 41], with the goal of obtaining a model that performs well in the source domain and generalizes well to unseen target domains by only providing a few samples to define how normal sounds belonging to these domains should sound like. Since DG is literally a generalization of DA to arbitrary domains rather than a specific target domain, it is more difficult to obtain such a model.
4.1 Datasets
Currently, there are four DCASE ASD datasets focusing on DG: the DCASE2022 [2], DCASE2023 [3], DCASE2024 [4], and DCASE2025 [5] datasets, which are all based on MIMII DG [42] as well as, respectively, ToyADMOS2 [14], ToyADMOS2+ [43], and ToyADMOS2# [44]. A part of the DCASE2024 dataset was recorded with the same setup as IMAD-DS [45]. IMAD-DS is another ASD dataset containing noisy multi-sensor signals of two machines in domain-shifted conditions recorded with a microphone, an accelerometer, and a gyroscope. The fundamental difference with the DCASE2021 dataset, which focuses on DA, is that the domain of individual test samples is unknown during inference. Moreover, the performance for each section is computed jointly for the source and target domains, i.e., a single decision threshold needs to be used for both domains. Other less severe differences include the size of all test sets, which is only half the size of the DCASE2021 test sets, and the availability of $10$ normal training samples belonging to the target domain instead of $3$. The differences between the DCASE2021 and DCASE2022 datasets end there, making them relatively similar. In contrast, the DCASE2023, DCASE2024, and DCASE2025 datasets only consist of a single section for each machine type, and the development and evaluation sets contain recordings of completely different machine types. Furthermore, the DCASE2024 and DCASE2025 datasets have noise conditions that are exclusively used for specific machine types, and for some machine types no additional attribute information is provided. The major difference between the DCASE2024 and DCASE2025 datasets is that supplemental recordings containing noise-free recordings of normal machine sounds or only background noise are provided for the train splits of the DCASE2025 dataset. Each year, modifications to the datasets result in a more realistic yet more challenging ASD task. For more details about these datasets, the reader is referred to Table 1 and the corresponding references.
4.2 Approaches
DG approaches for ASD in the context of DCASE can be grouped into four different strategies, which will now be discussed. Table 2 contains an overview of these approaches.
4.2.1 Domain Specialization
A simple approach for reducing domain mismatch is to balance the number of training samples belonging to the source and target domains. This can be achieved by balancing the domains in each mini-batch [21] or using more sophisticated approaches [22, 23, 24] such as synthetic minority over-sampling technique (SMOTE) [46]. However, since this approach trains models for specific domain shifts and thus requires to re-train the entire ASD system for each possible domain shift, this can be referred to as weak DG or domain specialization. Note that in contrast to DA, ASD systems need to work well for all domains without knowledge about the domain a given sample belongs to. Another way to handle this is to train domain-specific models and a domain classifier [21].
4.2.2 Domain-invariant Representations
The idea of domain-invariant representation learning [47] or domain-mixing-based approaches [2] is to reduce the variance between multiple domains based on the assumption that this also reduces the variance to other arbitrary target domains. In [25], individual samples of a batch are normalized independently of each other instead of using batch normalization [36] to avoid learning statistics that over-emphasize the source domain due to the highly imbalanced number of samples. The authors of [26] propose DG-mix, an extension of variance-invariance-covariance regularization (VICreg) [48] for self-supervised pre-training, using a loss term that minimizes the difference between domains and virtual domains created by mixup [49] before fine-tuning the model. Motivated by [50], which reduces the difference between second-order statistics of intermediate representations from domain-specific convolutional neural networks, [27] aims at reducing the difference between second-order statistics of source domain features and target domain features. To this end, the authors applied augmentations such as pitch shifting, time shifting, time stretching, adding white noise, and Filteraugment [51] to target domain samples to create more diverse domain shifts for training.
4.2.3 Feature Disentanglement
The main idea of feature disentanglement [52] is to decompose representations into domain-related features that are invariant for different domains and domain-unrelated features that capture the variability of different domains. The latter are independent of the domain and thus also generalize well to unseen domains. A possible way to achieve this is to use a domain-classification-based approach [2]. In most works on ASD in the DCASE context, this is achieved by focusing on the provided attribute information. The authors of [28] use two discriminative tasks for the sections and attributes. In [29], attribute information is disentangled in a normalizing-flow-based ASD model. In [30], a combination of a hierarchical metadata structure and attribute-specific Mahalanobis distances is used to learn more domain-related features. This approach is extended by [31] with gradient-reversal-based feature disentanglement of attribute information [53] and the use of a focal loss [54].
4.2.4 Anomaly Score Calculation
Last but not least, DG capabilities can also be improved by modifying the anomaly score computation. This approach has the advantage of not requiring expensive re-training of the neural networks that serve as the basis of state-of-the-art ASD systems. To this end, [32] uses an autoencoder and calculates the Mahalanobis distance between input data and reconstruction using domain-specific covariance matrices. For discriminative ASD systems, it was shown that simple nearest-neighbor-based anomaly scores lead to better results than estimating domain-specific distributions [33]. Further improvements can be obtained by normalizing the anomaly scores to reduce the domain mismatch. Examples are normalizing the anomaly scores based on local densities of normal reference samples [34] or by a domain-wise standardization of the anomaly score distributions [35].
5 Limitations and Future Directions
Table 3: Comparison of different types of DG approaches.
DG approach Effort Requires labels Effectiveness Domain Specialization high yes high Domain-invariant Representations medium no medium Feature Disentanglement low yes medium Anomaly Score Calculation very low sometimes low
All presented DG approaches have limitations. A comparison of the approaches can be found in Table 3. Although domain specialization approaches are highly effective, they can only adapt to specific domains and thus the entire ASD system, including the embedding model, needs to be re-trained for each domain shift, which is highly impractical as it is a strenuous process requiring access to domain labels. Obtaining domain-invariant representations does not require domain labels, but requires to simulate a diverse set of realistic domain shifts, which is a difficult task by itself and thus may limit the effectiveness. Feature disentanglement does not require much additional effort, but does require access to labeled data of various domains to be effective, which may not always be available. Focusing only on the calculation of the anomaly scores usually does not require one to modify the training of the embedding models. However, this approach is still sensitive to domain shifts if these also affect the obtained embeddings, thus limiting its effectiveness, even when making use of domain labels. Depending on the application and the available data, different DG approaches can be jointly used to mitigate the limitations of individual approaches. For example, anomaly score calculation approaches for DG can be used regardless of the learned representations, and simulated domain shifts, as used when learning domain-invariant representations, can also be used for feature disentanglement. However, the underlying issues of individual approaches persist and future work on DG is needed to solve these problems and identify potential synergies between specific approaches.
A future direction of research is continuous DA [55]. Here, the discrete indices for the domains are replaced with continuous ones to model the underlying relations between different domains. This improves consistency and may even allow to handle domains for which no training samples are available. Another direction is continual DA [56] that aims at handling domain shifts in non-stationary environments where domains are not static and may change continuously over time. In these environments, only adapting to a fixed domain is not sufficient. Note that continuous domain indices and continuously adapting to non-static domains can also be realized for DG. As a third direction, explaining the decisions of ASD systems is an open problem [57, 58], particularly in domain-shifted conditions. Here, future work may focus on explaining decisions of ASD in unknown target domains [59] or explaining domain shifts themselves [60]. Explainable ASD systems that focus on particular characteristics of the data may even offer better DG capabilities [61]. Last but not least, the semi-supervised ASD setting in the DCASE context can be relaxed to an unsupervised setting where even the training dataset may contain unlabeled anomalies [62].
6 Conclusion
In this work, we reviewed recent work on how to handle domain shifts for ASD tasks related to DCASE. We first defined domain shifts in the context of ASD, then discussed the topics of domain adaptation (DA) and domain generalization (DG) by motivating and defining these terms, presenting relevant datasets, and collecting different works related to these topics. In particular, DG approaches were grouped into domain specialization, domain-invariant representations, feature disentanglement, and anomaly score calculation. Furthermore, continuous and continual DA as well as explainable DG were identified as possible future research directions. For future work, we plan to compare all presented techniques in extensive experimental evaluations to identify the best working approaches and possible synergies.